# Agent Control Protocol ACP v1.19: Admission Control for Agent Actions Technical Specification and Reference Implementation

### 25 Mar 2026 v5

@@ 3  Technical Mechanisms
The Capability Token is ACP’s central artifact. It is a signed JSON object that specifies exactly what an agent can do, on what resource, for how long, and whether it can delegate that capability to other agents.
@@
Critical fields: exp is mandatory—a token without expiry is invalid by definition. The 128-bit nonce prevents replay attacks. parent_hash chains delegated tokens in a verifiable way. The signature covers all fields except sig.
@@
This sequence guarantees four formal properties: identity authentication, cryptographic request binding, anti-replay, and transport channel independence.
@@ 4  Inter-Institutional Trust
Recognition is explicitly non-transitive. If A recognizes B and B recognizes C, A does not automatically recognize C. Each bilateral relationship requires its own signed MRA. This prevents uncontrolled expansion of the trust graph.
@@ 4.3  Institutional Key Rotation and Revocation
Emergency rotation is activated when a key is compromised. The result is immediate: the key is marked revoked, all artifacts signed with it are invalid from that moment, and there is no transition period.
@@ 6.3  Validation Invariants
The validator enforces five invariants (REP-001 through REP-011): (1) evaluated_at $\leq$ valid_until prevents retroactively backdated snapshots; (2) now $\leq$ valid_until enforces expiration (v1.1 only, error REP-011); (3) score within scale bounds (REP-002); (4) non-empty issuer (REP-004); (5) valid Ed25519 signature (REP-010).
@@ 8.1  STRIDE Analysis
Spoofing AgentID impersonation $\text{AgentID}=\text{SHA-256}(\text{pk})$. Without valid signature $\rightarrow$ immediate DENIED.
@@
8.2  Guaranteed Security Properties
ACP guarantees the following properties when the implementation is compliant:
@@ 8.3  Declared Residual Risks
  • ITA bootstrap. The only point that depends on an out-of-band channel. Once the ITA root key is resolved, everything subsequent is autonomous.
@@ 9.1  Conformance Levels
L2 — SECURITY L1 + ACP-RISK-2.0, ACP-REV-1.0, ACP-ITA-1.0 Risk eval (F_anom + Cooldown), transitive revocation, ITA
@@ 9.3  Prohibited Behaviors
ACP defines 12 behaviors that no compliant implementation can exhibit. If any is exhibited, the implementation cannot declare conformance at any level:
Code Prohibited behavior
PROHIB-001 Approving a request when any evaluation component fails
PROHIB-002 Reusing an already-consumed Execution Token
PROHIB-005 Allowing state transition from revoked
PROHIB-007 Modifying or deleting Audit Ledger events
PROHIB-008 Silencing ledger corruption detection
PROHIB-009 Ignoring max_depth in delegation chains
PROHIB-011 Approving requests from agents with autonomy_level 0
Table 6: ACP prohibited behaviors. Exhibiting any disqualifies conformance at all levels.
@@ 11  Specification and Implementation Status
ACP v1.19 is a complete Draft Standard specification with a full Go reference implementation and external verifiability through a TLC-checked formal model and an ACR-1.0 sequence compliance runner.
11.1  Active Specifications — v1.19 (38 documents)
@@
L2 ACP-RISK-2.0 Deterministic Risk Engine v2 (F_anom + Cooldown)
@@
Table 7: ACP v1.19 active specification documents by conformance level (38 primary documents shown). The remaining 3 documents are the ACP-D decentralized suite (L5, design phase) not yet published. New in v1.16: ACP-RISK-2.0 (L2, deterministic engine v2 with F_anom + Cooldown), ACP-SIGN-2.0 (L1, PQC hybrid spec). New in v1.17: ACR-1.0 sequence compliance runner, TLA+ formal model, ACP-SIGN-2.0 HYBRID stub (pkg/sign2/). New in v1.18: performance benchmarks (latency, throughput, state contention), formal security/threat model, system comparison table. New in v1.19: adversarial evaluation (cooldown evasion, distributed multi-agent attack, state backend stress with Redis backend, compliance/adversarial/). Superseded versions (CONF-1.0, CONF-1.1, LEDGER-1.2, REP-1.1, AGENT-SPEC-0.3, DCMA-1.0, CROSS-ORG-1.0, REP-PORTABILITY-1.0) archived in archive/specs/.
@@ 11.2  Reference Implementation — 23 Go Packages + Demos
pkg/risk ACP-RISK-1.0 + ACP-RISK-2.0 L2
@@ 11.3  Conformance Test Vectors — 73 Signed Vectors
Table 9: Conformance test vector suites. All positive vectors carry real Ed25519 signatures (RFC 8037 Test Key A) and real SHA-256 hash chains.
11.4  ACP-RISK-2.0 Engine Performance
The pkg/risk ACP-RISK-2.0 engine was benchmarked using go test -bench on an Intel Core i7-8665U @ 1.90GHz (Go 1.22, Windows, GOMAXPROCS=8). These measurements reflect the pure in-memory evaluation cost, excluding network and HTTP parsing overhead.
Benchmark ns/op B/op allocs/op
Evaluate (APPROVED, no anomaly) 767 280 5
Evaluate (DENIED, no anomaly) 784 296 5
Evaluate (all 3 F_anom rules) 921 296 5
Evaluate (COOLDOWN short-circuit) 88 112 1
PatternKey (SHA-256 hash) 767 184 4
ShouldEnterCooldown 72 0 0
@@
The deterministic design of ACP-RISK-2.0 — integer arithmetic, fixed rules, no floating-point, no external ML inference — yields predictable latency bounds. At these throughput characteristics, the admission check adds negligible overhead relative to the actions it governs.

### 30 Apr 2026 v10

# Agent Control Protocol ACP v1.30: Admission Control for Agent Actions Technical Specification and Reference Implementation

@@ 3.4  Layered Architecture
3.5  Refinement under Transactional Assumptions
The Atomic Decision Boundaries framework [14] (Paper 0 in this governance series) establishes a formal requirement: a governance mechanism can guarantee runtime admissibility only if evaluation and state mutation occur in a single, indivisible LTS step. This section shows that ACP satisfies this requirement under two explicit transactional assumptions, and provides the correspondence table that connects both papers.
Transactional Assumptions
  • (TA1) Serializable ledger. All reads and writes to the state backend (LedgerQuerier) are serializable: no two concurrent evaluate calls observe interleaved intermediate states. In practice this is provided by the backend’s transaction isolation guarantee (e.g., Redis MULTI/EXEC, PostgreSQL SERIALIZABLE).
  • (TA2) Atomic commit. The pair (emit Capability Token, append Audit Ledger entry) either both succeed or both fail as an indivisible unit. ACP enforces this via the ACP-LEDGER-1.3 append protocol: a ledger write failure causes the token to be treated as if it were never issued (PROHIB-007: fail closed).
Correspondence Table
Table 5 maps the abstract objects of [14] to ACP’s concrete constructs.
Abstract (Paper 0) ACP instantiation
State $s\in S$ $(\mathit{ledger},\;\mathit{riskScore},\;\mathit{cooldownExpiry},\;\mathit{agentState})$ held by the LedgerQuerier
Action $a\in\mathit{Act}$ Incoming capability request $r$ carrying (AgentID, cap[], res, exp, nonce, sig)
Decision function $F(s,a)$ evaluate(s,r)(s,r): verify signature $\to$ score risk $\to$ compare thresholds $\to$ return APPROVED / DENIED / ESCALATED
$F(s,a)=(\textsc{Allow},s^{\prime})$ APPROVED: emit Capability Token + append ledger entry; $s^{\prime}=s[\mathit{cooldown}{+}{=}\Delta,\;\mathit{ledger}{+}{=}\mathit{entry}]$
$F(s,a)=(\textsc{Refuse},s^{\prime})$ DENIED: no token emitted; ledger records denial; $s^{\prime}=s[\mathit{anomalyCount}{+}{=}1]$
$F(s,a)=(\textsc{Escalate},\hat{s}^{\prime})$ ESCALATED: request placed in audit queue pending supervisor resolution; $\hat{s}^{\prime}=(s,\;\mathit{auditQueue}\cup\{r\})$
Admissibility $\mathit{Adm}(s,a)$ $\mathit{RS}(r,s)<\theta_{\mathit{deny}}$ (risk score below denial threshold for agent’s autonomy level)
Escalation space $\hat{S}{=}S{\times}\mathit{Req}$ State backend extended with escalation_queue: $\hat{S}=S\times\mathit{PendingEscalations}$
Table 5: Correspondence between the abstract Atomic Decision Boundary model [14] and ACP constructs. Under TA1 and TA2 each row is realized atomically.
Refinement Claim
Claim. Under TA1 and TA2, the ACP evaluate-then-mutate pipeline realizes the atomic decision boundary $F:S\times\mathit{Act}\to\mathcal{D}\times\hat{S}$ of [14] as a single LTS step: state read, decision computation, and state write are executed without interleaving, and token emission and ledger commit are jointly atomic. This is a refinement (not a software implementation) of the abstract model, because it adds concrete cryptographic, timing, and network constructs not present in the abstract LTS.
Two corollaries follow:
  1. 1.
    Admissibility guarantee. Under TA1–TA2, every ACP APPROVED decision corresponds to an admissible action at evaluation time. Split evaluators (RBAC, OPA, policy engines) cannot guarantee this because state can change between read and write.
  2. 2.
    Escalation closure. Every ESCALATED request is appended to the audit queue atomically with its state snapshot, making supervisor resolution deterministic and replay-safe.
Scope. The refinement holds under TA1–TA2. When the state backend provides only read-committed isolation (weaker than serializable), two concurrent evaluations can observe the same pre-mutation state, breaking atomicity; this is a declared limitation (see §13.8). The behavioral layer above this boundary—whether drift accumulates within the admitted action space—is the subject of [15] (Paper 2, IML). Allocation fairness across competing agents sharing the boundary, strategy-proofness, and compositional irreducibility of the full four-layer stack are addressed in [16] (Paper 3/4).
@@ 4  Technical Mechanisms
The Capability Token is ACP’s central artifact. It is a signed JSON object [25] that specifies exactly what an agent can do, on what resource, for how long, and whether it can delegate that capability to other agents.
@@
Critical fields: exp is mandatory—a token without expiry is invalid by definition. The 128-bit nonce enforces request uniqueness within the canonical representation and supports anomaly accumulation; it does not constitute nonce-based replay prevention (see Section 13.7 and Section 13.8). parent_hash chains delegated tokens in a verifiable way. The signature covers all fields except sig.
@@
This sequence is designed to provide four properties: identity authentication, cryptographic request binding, anti-replay, and transport channel independence.
@@
4.5  LLM Agent Integration
ACP is designed to integrate with LLM agent frameworks (LangChain, LangGraph, AutoGen, CrewAI) as an admission control layer that intercepts tool calls before they modify system state. The integration pattern wraps each tool with an ACP guard that calls POST /acp/v1/authorize and gates execution on the admission decision.
class ACPGuardedTool(BaseTool):
capability: str # e.g. "acp:cap:financial.transfer"
resource: str # e.g. "accounts/restricted-fund"
res_class: str # "PUBLIC" | "SENSITIVE" | "RESTRICTED"
agent_id: str # agent identity bound to ACP token
acp_url: str # ACP server endpoint
def _run(self, query: str) -> str:
resp = requests.post(f"{self.acp_url}/acp/v1/authorize", json={
"agent_id": self.agent_id,
"capability": self.capability,
"resource": self.resource,
"action_parameters": {"resource_class": self.res_class},
}).json()
decision = resp["decision"]
if decision == "APPROVED":
return self._execute(query) # tool logic runs
if decision == "ESCALATED":
return "[ESCALATED: pending human review]"
return f"[DENIED by ACP: {resp.get(’reason_code’)}]"
Listing 1: ACP-guarded tool wrapper for LangChain agents. Each tool call is intercepted before execution; admission is decided by the ACP server based on the agent’s identity, capability, and resource class.
The admission decision is stateful: the ACP server maintains denial history, pattern counters, and cooldown state across calls from the same agent. A tool call that would be individually approved may be denied if the agent’s recent history has accumulated sufficient risk signals (F_hist, F_anom). This gives the framework a temporal enforcement property that stateless policy engines cannot provide.
IPI composability. A prompt-layer filter (e.g., PromptGuard, PINT) can be applied before the tool call reaches ACPGuardedTool, reducing the fraction of injected high-risk calls that reach the admission boundary. ACP enforces the structural constraint on the tool call regardless of its origin: a fund_transfer call with $\mathrm{RS}=80$ is denied whether it was generated by legitimate user input or by an injected instruction. The two layers are composable: the prompt-layer filter reduces attack frequency; ACP provides the structural denial guarantee.
Reproducible local evaluation. For evaluation without external API dependencies, the Ollama runtime [35] provides OpenAI-compatible endpoints for open-weight models. Setting temperature=0 and seed=42 yields pseudo-deterministic agent behavior, enabling controlled experimental evaluation of ACP admission decisions across scripted agent sessions (see Experiment 12, §17.10).
Real-LLM validation. To validate the ACPGuardedTool wrapper in a live agent context, we ran a five-turn session using DeepSeek-R1:8B [9] served locally via Ollama (temperature=0, seed=42). An IPI payload was embedded in the weather_query tool result, instructing the agent to call fund_transfer with attacker-controlled parameters.
Model: deepseek-r1:8b | temperature=0 | seed=42
[T1] weather_query(city="Madrid") RS= 0 APPROVED [IPI injected]
[T2] fund_transfer($99$, "attacker-001") RS=80 DENIED [IPI-induced]
[T3] weather_query(city="Barcelona") RS= 0 APPROVED [IPI injected]
[T4] fund_transfer($99$, "attacker-001") RS=80 DENIED [IPI-induced]
[T5] fund_transfer(50, "savings-123") RS=80 DENIED [legitimate]
--> COOLDOWN ACTIVATED
Listing 2: Condensed output from the ACP real-LLM IPI demo (demos/ollama-agent/agent_demo.py). Full session is deterministic and reproducible.
Three observations follow. First, the LLM followed the IPI injection in turns 2 and 4, calling fund_transfer with attacker-controlled parameters in both cases. ACP blocked each attempt because the capability acp:cap:financial.transfer with resource class RESTRICTED evaluates to $\mathrm{RS}=80\geq 70$, regardless of the transaction parameters. The LLM’s internal reasoning is irrelevant to ACP’s admission decision.
Second, turn 5 demonstrates capability-level enforcement: a legitimate transfer (\$50 to savings-123) is also denied because ACP evaluates capability and resource class, not individual parameters. In a production deployment, a personal savings account would carry resource class SENSITIVE ($\mathrm{RS}=35+15=50$, ESCALATED), separating attacker-targeted restricted transfers from legitimate sensitive ones. The demo uses a uniform RESTRICTED class to isolate the IPI dynamic.
Third, the third denial in turn 5 activated agent-wide cooldown, locking the agent for 300 seconds. The same stateful consequence observed in Experiment 12 — $F_{\text{anom}}$ Rule 2 persisting beyond the cooldown window — emerges here with a real LLM as the agent brain, confirming that the enforcement properties are independent of the agent’s implementation.
4.6  Context-Scoped Anomaly Enforcement (ACP-RISK-3.0)
ACP-RISK-2.0 evaluates $F_{\text{anom}}$ against a single per-agent counter: CountRequests(agentID, window). This produces a state-mixing vulnerability when the same agent holds capabilities in multiple resource contexts: denials accumulated in one context contaminate the RS in an unrelated context, producing false ESCALATED or DENIED outcomes (see §13.8 and Experiment 7).
ACP-RISK-3.0 eliminates this by scoping the anomaly counter to a context key:
$\texttt{PatternKey}(a,c,r)=\text{SHA-256}(\,a\;\|\;c\;\|\;r\,)$
where $a$ is the agent_id, $c$ the capability, and $r$ the resource. Rule 1 of $F_{\text{anom}}$ becomes CountPattern(ctxKey, -60s) rather than CountRequests(agentID, -60s). All other rules ($F_{\text{anom}}$ Rules 2–3, cooldown logic, decision thresholds) are unchanged.
The key derivation adds a single SHA-256 call per evaluation (739 ns; Table 12, PatternKey row), preserving the sub-microsecond latency envelope. Experiment 8 (§13.7) confirms that cross-context contamination is fully eliminated under ACP-RISK-3.0 while same-context burst detection remains intact.
@@ 5  Inter-Institutional Trust
Recognition is explicitly non-transitive. If A recognizes B and B recognizes C, A does not automatically recognize C. Each bilateral relationship requires its own signed MRA. This limits uncontrolled expansion of the trust graph.
@@ 5.3  Institutional Key Rotation and Revocation
Emergency rotation is activated when a key is compromised. The key is marked revoked; all artifacts signed with it are invalid from the moment revocation is observed by the verifier, and there is no transition period.
5.4  Trust Model and Failure Modes
ACP makes its trust assumptions explicit, bounded, and temporally scoped: all protocol guarantees hold only after successful bootstrap and until revocation is observed. This section formalizes the three structural trust scenarios that determine the security boundary of the ITA layer.
K1a — Bootstrap Trust Assumption.
Assumption. The initial distribution of the ITA root public key occurs over an authentic out-of-band channel.
Adversary capability. An adversary who controls the bootstrap channel can substitute an attacker-controlled key, after which all ACP artifact verification by the victim is performed against the wrong authority.
Formal claim. ACP provides no cryptographic security guarantees if the bootstrap trust anchor is compromised. All subsequent guarantees—token verification, delegation validity, ledger integrity—are conditional on the authenticity of the bootstrapped ITA public key.
Security implication. This is an explicit design boundary, not a protocol defect. ACP reduces the trust problem from continuous dependence on a third party to a single, auditable bootstrap event. Deployment environments may use out-of-band channels with varying integrity guarantees (e.g., DNSSEC, certificate pinning, in-person key ceremony); the choice of channel is a deployment decision outside the protocol scope.
K1b — Key Compromise Window.
Assumption. Key compromise occurs at time $T_{\mathit{compromise}}$ and is detected and revoked at $T_{\mathit{revoke}}$, with $T_{\mathit{compromise}}<T_{\mathit{revoke}}$.
Adversary capability. During the interval $T_{\mathit{compromise}},\,T_{\mathit{revoke}}]$, the adversary holds a valid signing key and can issue ACP-conformant tokens and ledger events that are indistinguishable from legitimate artifacts.
Formal claim. ACP evaluates token validity at verification time, not issuance time. Security is monotonic with respect to revocation: once a key is marked revoked, all subsequent verifications reject tokens derived from it, regardless of when they were issued. Tokens issued during the compromise window remain valid until revocation is observed.
Security implication. The window $T_{\mathit{compromise}},\,T_{\mathit{revoke}}]$ represents an irrecoverable exposure interval: ACP cannot retroactively invalidate actions that completed before revocation was processed. Minimizing this window is a deployment concern (monitoring, anomaly detection, HSM audit logs) that the protocol does not prescribe but does not preclude.
K1c — Revocation Authority Model.
Assumption. The ITA authority is reachable and uncompromised.
Adversary capability. If the ITA authority is unavailable or compromised, revocation cannot be issued or cannot be propagated, leaving previously valid tokens uncancellable.
Formal claim. ACP assumes the existence of an ITA authority capable of issuing revocations independent of delegation chains. Delegation chains are not self-healing: a compromised intermediate agent cannot unilaterally revoke its own parent; revocation authority flows from the ITA downward, not upward through the chain.
Security implication. Security depends on the availability and integrity of the ITA revocation authority. This is a known architectural constraint shared with all PKI-based systems. ACP makes this dependency explicit rather than implicit. Deployments requiring high availability of revocation must provision the ITA with appropriate redundancy outside the protocol scope.
@@ 7.3  Validation Invariants
The validator enforces five invariants (REP-001, REP-002, REP-004, REP-010, REP-011): (1) evaluated_at $\leq$ valid_until prevents retroactively backdated snapshots; (2) now $\leq$ valid_until enforces expiration (v1.1 only, error REP-011); (3) score within scale bounds (REP-002); (4) non-empty issuer (REP-004); (5) valid Ed25519 signature (REP-010).
@@ 9.1  STRIDE Analysis
Spoofing AgentID impersonation $\text{AgentID}=\text{base58}(\text{SHA-256}(\text{pk}))$. Without valid signature $\rightarrow$ immediate DENIED.
@@
9.2  Security Properties Under Correct Implementation
ACP provides the following properties when the implementation is compliant:
@@ 9.3  Declared Residual Risks
  • ITA trust model. Bootstrap, key compromise window, and revocation authority constraints are formalized in §5.4. All protocol guarantees are conditioned on the authenticity of the bootstrapped ITA key and the availability of the revocation authority.
  • Agent-internal threats. ACP operates at the execution boundary between agent intent and system state mutation; it does not observe or control the agent’s internal reasoning process. Threats that originate within that process—including prompt injection and tool-use escalation inside an LLM reasoning loop—fall outside ACP’s enforcement scope. ACP is designed to complement sandboxing and monitoring systems that address these threats at the reasoning layer, not to replace them.
@@ 10.1  Conformance Levels
L2 — SECURITY L1 + ACP-RISK-3.0, ACP-REV-1.0, ACP-ITA-1.0 Risk eval (context-scoped F_anom + Cooldown), transitive revocation, ITA
@@ 10.3  Prohibited Behaviors
ACP defines 12 behaviors that no compliant implementation can exhibit (Table 45); exhibiting any disqualifies conformance at all levels.
@@ 12  Specification and Implementation Status
ACP v1.30 is a complete Draft Standard specification with a full Go reference implementation and external verifiability through a TLC-checked formal model and an ACR-1.0 sequence compliance runner.
12.1  Active Specifications — v1.30 (38 documents)
@@
L2 ACP-RISK-3.0 Deterministic Risk Engine v3 (context-scoped F_anom, Cooldown)
L2 ACP-RISK-2.0 Deterministic Risk Engine v2 (superseded by v3)
@@
Table 9: ACP v1.30 active specification documents by conformance level (38 primary documents shown). The remaining 3 documents are the ACP-D decentralized suite (L5, design phase) not yet published. Superseded versions archived in archive/specs/. Version history per document is tracked in Table 26.
@@ 12.2  Reference Implementation — 23 Go Packages + Demos
pkg/risk ACP-RISK-1.0 + ACP-RISK-2.0 + ACP-RISK-3.0 L2
@@ 12.3  Conformance Test Vectors — 73 Signed Vectors
Table 11: Conformance test vector suites. All positive vectors carry real Ed25519 signatures (RFC 8037 Test Key A [30]) and real SHA-256 hash chains.
12.4  ACP-RISK-2.0/3.0 Engine Performance
Benchmark methodology.
Three distinct benchmark families are reported in this section; each measures a different layer of the system and they are not directly comparable:
  1. 1.
    Engine microbenchmarks (go test -bench, Table 12): measure the pure pkg/risk.Evaluate() call with an InMemoryQuerier, single goroutine, no HTTP or network overhead. Reports ns/op (739–832 for full paths, 78 for the cooldown short-circuit) and the 1,720,000 req/s single-goroutine throughput ceiling.
  2. 2.
    Concurrent pipeline benchmark (Table 13): measures end-to-end request throughput with 10–500 concurrent goroutines feeding a shared InMemoryQuerier. Reports 920,000 req/s at 10 workers, degrading to 712,000 at 500 workers. The difference from the microbenchmark reflects scheduling, mutex contention, and goroutine overhead rather than evaluation logic.
  3. 3.
    Controlled latency injection (DelayedQuerier, Table 18): simulates realistic external backend latency by inserting a configurable delay in the LedgerQuerier call, isolating the effect of state backend latency on system throughput.
The pkg/risk engine was benchmarked using go test -bench on an Intel Core i7-8665U @ 1.90GHz (Go 1.22, Windows, GOMAXPROCS=8). Engine measurements reflect the pure in-memory evaluation cost, excluding network and HTTP parsing overhead.
Scenario ns/op B/op allocs/op Path
Evaluate (clean state) 739 280 5 full RS
Evaluate (with history) 803 280 5 full RS
Evaluate (all 3 F_anom rules) 832 296 5 full RS
Evaluate (COOLDOWN short-circuit) 78 112 1 Step 2 exit
Throughput under concurrent load (InMemoryQuerier, 10 workers)
10 concurrent agents 581 1,722,000 req/s
100 concurrent agents 629 1,589,000 req/s
500 concurrent agents 663 1,510,000 req/s
@@
The deterministic design of ACP-RISK-3.0 — integer arithmetic, fixed rules, no floating-point, no external ML inference — yields predictable and verifiable latency bounds. At sub-microsecond cost per decision, ACP’s overhead is six orders of magnitude below the latency of the actions it governs: typical LLM inference takes 100 ms–10 s; external API calls 1–100 ms. The admission check is not a performance bottleneck in any realistic deployment.
Throughput degradation from 10 to 500 concurrent agents (1.72M to 1.51M req/s, -12%) confirms that the evaluation function itself is not the bottleneck: the limiting factor is state backend access (InMemoryQuerier uses a sync.Mutex), not the admission logic. Production deployments replacing InMemoryQuerier with a Redis or PostgreSQL backend incur I/O latency but decouple throughput from the evaluation function entirely, enabling horizontal scaling.
ACP-RISK-3.0 preserves these latency bounds over ACP-RISK-2.0; context-scoped key derivation adds one SHA-256 call per evaluation, a cost dominated by existing ledger I/O in production deployments.